Calculate the one temperature at which Fahrenheit and Celsius thermometers agree with each other

1
Expert's answer
2022-01-24T18:43:55-0500

Formula for Fahrenheit,"F = \frac{9}{5 C + 32}"


Formula for Centigrade, "C =\frac{5}{9 (F-32)}"


Let F = C or C= F


"F= C=\frac{ 5}{9(F-32)}=\frac{ (5F-160)}{9}"

9F= 5F - 160


9F - 5F = -160


4F = -160


F = "-\frac{160}{4}"


F= -40 degree answer in Fahrenheit


C= F= "\frac{9}{5 C + 32}"


C = "\frac{(9C + 160)}{5}"


5C = 9C + 160


5C - 9C = 160


-4C= 160


C="\frac{160}{-4}"


C= -40 degree answer in Centigrade


Conclusion: The temperature wherein degrees centigrade is the same in degrees Fahrenheit is -40 degrees .